1. Hydrogen sulfide decomposes according to the following reaction, for which Kc = 0.93 at 700 °C:


2H2S(g) <---> 2H2(g) + S2(g)


If 0.75 mol of H2S is placed in a 3.0-L container, what are the equilibrium concentrations of all the species?


1
Expert's answer
2021-10-30T01:15:07-0400

The balanced reaction equation is: 2H2S(g)⇋2H2(g)+S2(g)

The equilibrium constant expression is: Kc=9.30×10-8

="\frac{2H2×S2}{H2S}"

Initially, we only have the reactant:

"[\nH\n2\nS\n]\n=\n0.45\n \nm\no\nl\n\u00f7\n3.0\n \nL\n=\n0.15\n \nM"

"[\nH\n2\nS\n]\ne\nq\n\u2248\n0.15\n \nM"

"H\n2\n]\ne\nq\n=\n2\nx"

"[\nS\n2\n]\ne\nq\n=\nx"

"9.30
×
10
^{−
8}
=
\frac{(
2
x
)^
2
(
x
)}{
(
0.15
 
M
)
^2}"

"[H\n2]\n\ne\nq\n=\n2\n\u00d7\n8.06\n\u00d7\n10^{\n\u2212\n4}M"

"=\n1.61\n\u00d7\n10\n^{\u2212\n3}\n \nM"